The vapor pressure of liquid bromoethane, C2H5Br, is 100. mm Hg at 263 K. A sample of C2H5Br is placed in a closed, evacuated 537 mL container at a temperature of 263 K. It is found that all of the C2H5Br is in the vapor phase and that the pressure is 57.0 mm Hg. If the volume of the container is reduced to 328 mL at constant temperature, which of the following statements are correct?

Choose all that apply.

No condensation will occur.

Liquid bromoethane will be present.

Only bromoethane vapor will be present.

Some of the vapor initially present will condense.

The pressure in the container will be 100. mm Hg.

Please list from above which would apply to this problem thank you!
Answer #1

Correct statements are :

(1) No condensation will occur.

(2) Only bromoethane vapor will be present.

Explanation

initial pressure = 57.0 mmHg

initial volume = 537 mL

final volume = 328 mL

According to Boyle's law,

P1 * V1 = P2 * V2

P1 = initial pressure = 57.0 mmHg

V1 = initial volume = 537 mL

P2 = final pressure

V2 = final volume = 328 mL

P2 = P1 * (V1 / V2)

P2 = 57.0 mmHg * (537 mL / 328 mL)

P2 = 57.0 mmHg * (1.637)

P2 = 93.3 mmHg

Since this pressure is less than vapor pressure, no condensation will occur and only vapor will be present.